Luis weighs 54 more than carla. Carla weighs 2 pounds more than rita. Together their weights total 333 pounds. How much does each weigh?
Answer by mananth(16949)

Rita be x pounds
Carla = x+2
Luis = x+2+54
...
x+x+2+x+2+54=333
3x+58=333
add -58 to both sides
3x+58-58=333-58
3x=275
divide by 3
x= 91.66 pounds. Rita
..
Carla = 91.66+2 = 93.66 pounds
Luis = 93.66+54=147.66
..
CHECK
91.66+93.66+147.66=322.98